# Spectral decomposition of field operators and causal measurement in quantum field theory

### v1

@@ Appendix A Appendix
$\int_{-\infty}^{\infty}\mathrm{d}s\,\langle\mathsf{K}_{\gamma},\Pi_{\xi^{\prime}}^{\epsilon^{\prime}}(q-s)\Pi_{\xi}^{\epsilon}(s)\mathsf{K}_{\beta}\rangle=\int_{-\infty}^{\infty}\mathrm{d}s\,\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\pi\sqrt{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(2\{\xi,\xi^{\prime}\}\left(D_{\xi}(\beta,\gamma)-s\right)\left(D_{\xi^{\prime}}(\beta,\gamma)-(q-s)\right)\right.\right.\\ \left.\left.-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)-s\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-(q-s)\right)^{2}\right)\right)\\ =\int_{-\infty}^{\infty}\mathrm{d}s\,\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\pi\sqrt{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(-\left(2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}\right)s^{2}\right.\right.\\ +2\left((\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})D_{\xi}(\beta,\gamma)-(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})(D_{\xi^{\prime}}(\beta,\gamma)-q)\right)s\\ \left.\left.+2\{\xi,\xi^{\prime}\}D_{\xi}(\beta,\gamma)\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)^{2}\right)\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\sqrt{\pi}\sqrt{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(\frac{1}{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}\right.\right.\\ \left.\left((\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})D_{\xi}(\beta,\gamma)-(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})(D_{\xi^{\prime}}(\beta,\gamma)-q)\right)^{2}\right)\\ \left.+2\{\xi,\xi^{\prime}\}D_{\xi}(\beta,\gamma)\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)^{2}\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\sqrt{\pi}\sqrt{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\frac{1}{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}\right.\\ \left((\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})^{2}(D_{\xi}(\beta,\gamma))^{2}+(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})^{2}(D_{\xi^{\prime}}(\beta,\gamma)-q)^{2}\right.\\ \left.-2(\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})D_{\xi}(\beta,\gamma)(D_{\xi^{\prime}}(\beta,\gamma)-q)\right.\\ +\left(2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}\right)\\ \left.\left.\left(2\{\xi,\xi^{\prime}\}D_{\xi}(\beta,\gamma)\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)^{2}\right)\right)\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\sqrt{\pi}\sqrt{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}}\\ \exp\left(-\frac{1}{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}\left(D_{\xi}(\beta,\gamma)+D_{\xi^{\prime}}(\beta,\gamma)-q\right)\right)$ (93)

### v2

@@ Appendix A Appendix
$\int_{-\infty}^{\infty}\mathrm{d}s\left(B_{\xi^{\prime}}^{\epsilon^{\prime}}(q-s)\star B_{\xi}^{\epsilon}(s)\right)(\mathsf{K}_{\gamma},\mathsf{K}_{\beta})=\int_{-\infty}^{\infty}\mathrm{d}s\,\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\pi\sqrt{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(2\{\xi,\xi^{\prime}\}\left(D_{\xi}(\beta,\gamma)-s\right)\left(D_{\xi^{\prime}}(\beta,\gamma)-(q-s)\right)\right.\right.\\ \left.\left.-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)-s\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-(q-s)\right)^{2}\right)\right)\\ =\int_{-\infty}^{\infty}\mathrm{d}s\,\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\pi\sqrt{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(-\left(2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}\right)s^{2}\right.\right.\\ +2\left((\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})D_{\xi}(\beta,\gamma)-(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})(D_{\xi^{\prime}}(\beta,\gamma)-q)\right)s\\ \left.\left.+2\{\xi,\xi^{\prime}\}D_{\xi}(\beta,\gamma)\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)^{2}\right)\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\sqrt{\pi}\sqrt{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(\frac{1}{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}\right.\right.\\ \left.\left((\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})D_{\xi}(\beta,\gamma)-(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})(D_{\xi^{\prime}}(\beta,\gamma)-q)\right)^{2}\right)\\ \left.+2\{\xi,\xi^{\prime}\}D_{\xi}(\beta,\gamma)\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)^{2}\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\sqrt{\pi}\sqrt{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\frac{1}{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}\right.\\ \left((\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})^{2}(D_{\xi}(\beta,\gamma))^{2}+(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})^{2}(D_{\xi^{\prime}}(\beta,\gamma)-q)^{2}\right.\\ \left.-2(\{\xi,\xi^{\prime}\}+\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})(\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\epsilon^{2})D_{\xi}(\beta,\gamma)(D_{\xi^{\prime}}(\beta,\gamma)-q)\right.\\ +\left(2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}\right)\\ \left.\left.\left(2\{\xi,\xi^{\prime}\}D_{\xi}(\beta,\gamma)\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q\right)^{2}\right)\right)\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\sqrt{\pi}\sqrt{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}}\\ \exp\left(-\frac{1}{2\{\xi,\xi^{\prime}\}+\|\xi\|^{2}+\|\xi^{\prime}\|^{2}+\epsilon^{2}+\epsilon^{\prime 2}}\left(D_{\xi}(\beta,\gamma)+D_{\xi^{\prime}}(\beta,\gamma)-q\right)\right)$ (94)